the average of 5 numbers is 10. the first four numbers are 8, 9, 10, 11. find the fifth number
Found 2 solutions by FrankM, stanbon:
Answer by FrankM(1040)

For 5 number to average 10, they must add to 50, as 50/5 = 10, right?
Your 4 numbers sum to 38, so the missing number is 50 - 38 = 12

Answer by stanbon(75887)

the average of 5 numbers is 10. the first four numbers are 8, 9, 10, 11. find the fifth number
-----
Let the 5th number be "x":
------
Equation:
(8+9+10+11+x)/5 = 10
-----
(38+x) = 50
x = 12
===============
Cheers,
Stan H.